a particle executing a simple harmonic motion has a period of 6s the time period taken by the particle to move from the mean position to half the amplitude starting from the mean position is
 

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
X = A sin wt
At x = A/2
Sin wt = ½
wt = pi/6
Now
t = pi/6*w
t = 6T/2*pi = T/12
T/12 = 6/12 = 1/2 sec
 
Regards
Arun (askIITians forum expert)